Circular No. 7659

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

X-RAY TRANSIENTS IN M31 R. Shirey, University of California at Santa Barbara, on behalf of the

X-Ray Multi-Mirror (XMM) Newton Optical/Ultraviolet Monitor and the XMM-Newton Survey Science Centre teams, reports the discovery of two new transient x-ray sources in M31: “The transients were detected in ∼ 50 000-s exposures with each of the three EPIC instruments during XMM-Newton observations centered on June 29.52 UT. One of the transients (XMMU J004308.5+411820), located at α = 0^{h}43^{m}08^{s}.5, δ = +41^{o}18′20′′ (equinox 2000.0; ∼ 3′′ error radius, based on astrometry relative to Chandra globular-cluster x-ray source positions), exhibits a supersoft x-ray spectrum (thermal spectrum with kT ∼ 30 eV) with an observed flux (0.2–2 keV) of 3 × 10^{−14} erg s^{−1} cm^{−2}. During the x-ray observations, the Optical/ultraviolet Monitor (OM) onboard XMM obtained exposures in three ultraviolet filters. A faint ultraviolet source is present within 2′′–3′′ of the location of the supersoft x-ray transient with fluxes of 6 × 10^{−17}, 5 × 10^{−17}, and 6 × 10^{−17} erg s^{−1} cm^{−2} Å^{−1} in UVW1 (245–320 nm), UWM2 (205–245 nm), and UVW2 (180–225 nm) exposures of 5000, 15 000, and 21 000 s, respectively. Shorter B-band and UVW1 OM exposures (800 and 1200 s) from Dec. 2000 show no such source at this location. In the MIT/Amsterdam M31 Survey, no source is catalogued down to B ∼ 21–22 within 5′′ of the transient location, while several stars of magnitude I ∼ 19 lie within 4′′ of the transient. A second new transient x-ray source (XMMU J004245.2+411721) is located at α = 0^{h}42^{m}45^{s}.2, δ = +41^{o}17′21′′ (equinox 2000.0; ∼ 3′′ error radius) and exhibits a more typical power-law spectrum (photon index ∼ 1.5) and observed flux (0.2–10 keV) 1.4 × 10^{−13} erg s^{−1} cm^{−2}. Finally, the near-nuclear x-ray transient discovered with Chandra in Oct. 1999 (IAUC 7291) and seen to be ‘on’ by XMM-Newton in June and Dec. 2000 is no longer detected. Observations of the two new x-ray transients at other wavelengths are encouraged.”

COMET C/2001 A2 (LINEAR) Total visual magnitude estimates: July 6.05 UT, 5.0 (J. Kysely, Vlasim,

Czech Republic, 7×50 binoculars); 7.98, 5.0 (S. Baroni, Milan, Italy, 40×80 binoculars; moonlight); 10.89, 5.1 (K. Hornoch, Lelekovice, Czech Republic, 10×80 binoculars); 12.41, 5.1 (C. S. Morris, Fillmore, CA, 10×50 binoculars; two tails, each at least 2^{o} long).

2001 July 12  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green